# Conditional score-based diffusion models for solving inverse problems in mechanics

### v2

@@ 1 Introduction
Clearly, the conditional generative modeling resembles the process of solving an inverse problem within the Bayesian paradigm where ${\mathbf{X}}$ and ${\mathbf{Y}}$ now represent the quantity that must be inferred and the measurements available, respectively [17]. Thus, motivated by the recent wide-scale adoption of diffusion models, we proposed to use conditional diffusion models to solve a class of inverse problems in mechanics that involve inferring the spatially varying material constitutive parameters from full-field measurements of the mechanical response, such as displacements or strains. Inverse problems of this type have many applications in material characterization [18], structural health monitoring and damage detection [19, 20], and medical diagnosis [21, 22]. We are particularly motivated by the applications of this type of inverse problem in medical diagnosis, which arises when using elastography-based imaging techniques that assess the state of tissue using its mechanical properties, mainly to differentiate between healthy and diseased tissue [21, 23].
@@
Solving inverse problems of the aforementioned type can be challenging. First, the inferred quantity is a field that must be discretized to make the inverse problem tractable, leading to a high-dimensional inverse problem. Second, the forward model is complicated. Mechanistic models that simulate the deformations of the specimen in consideration may be a high-fidelity black-box model. Moreover, the measurement noise in these applications may be complex— non-Gaussian and non-additive. These factors combine to make Bayesian inference difficult since standard inference tools, such as Markov chain Monte Carlo (MCMC) methods and variational inference, may not perform well in the face of these challenges. However, the recent success of conditional diffusion models in solving high-dimensional inverse problems offers promise [24, 25, 26, 27, 28].
@@
We propose to use conditional score-based diffusion models — a particular type of diffusion model that leverages deep neural networks to approximate the score function of the target posterior distribution [29, 30, 31]. In the context of inverse problems, the score function is the logarithm of the posterior density. The neural network approximating the score function is called the score network [31]. Approximating the score function allows greater flexibility in choosing the architecture for the score network in the absence of additional constraints necessary to model a valid probability density. For instance, the score network need not be invertible [31], unlike normalizing flows [32]. Moreover, a single neural network can approximate the posterior distribution for different realizations of the covariate ${\mathbf{Y}}$, which only requires adding a channel to the input of the neural network [33, 34, 35]. This also allows for re-using the trained conditional score-based diffusion for different realizations of ${\mathbf{Y}}$ during testing. Significantly, we show that the score network can be trained using a regression-type objective, see Eq. 15, which is much more stable than adversarial losses used to train GANs. Moreover, training only requires samples from the joint distribution between ${\mathbf{X}}$ and ${\mathbf{Y}}$. The training dataset can be easily created by simulating the forward model for different realizations of ${\mathbf{X}}$ sampled from a suitable prior or collected a priori through experiments. Thus, the forward model can be arbitrarily complex, black-box, and incompatible with automatic differentiation. This lends great utility to the proposed approach since most forward models for the inverse problems of practical interest are likely to have one or more of the aforementioned properties. Moreover, the training dataset is entirely synthetic and does not require access to experimental data, which can be scarce in some applications. After training the score network, we can sample the posterior distribution for any realization of ${\mathbf{Y}}$, perhaps the result of an experiment, using Langevin dynamics MCMC [31, 36].
@@
We note that recent work by Jacobsen et al. [37] uses diffusion models for full-field inversion from sparse measurements. Jacobsen et al. [37] primarily focuses on enforcing physical consistency to the realizations generated using the conditional diffusion model, which is not the focus of the current work. The realizations obtained using conditional score-based diffusion models can be enhanced by enforcing physical consistency using ideas developed in [37]. Similarly, Bastek et al. [38] propose to enforce physical consistency during training of the score network, which can also be used to augment the proposed approach. However, enforcing physical consistency by discretizing the underlying forward model may not be straightforward in many applications involving nonlinear phenomena and complex geometries. Therefore, we do not pursue informing the score network or the sampling process with physics knowledge in this work. Indeed, the proposed approach is agnostic to the physics of the forward process, i.e., we assume that the forward model is black-box.
@@
The remainder of this paper is organized as follows. In Section 2, we set up the inverse problem we intend to solve and provide the necessary background on score-based diffusion models. We introduce our proposed approach to solve inverse problems using conditional score-based diffusion models in Section 3 where Section 3.1 develops the training objective, Section 3.2 outlines the sampling scheme, Section 3.3 provides guidelines to choose various hyper-parameters associated with the model and sampling, Section 3.4 provides design details pertaining to the neural network used to approximate the score function, and Section 3.5 provides an overview of the various steps in the proposed approach. In Section 4, we use the proposed approach to solve various inverse problems that concern the inference of a spatially varying constitutive parameter from noisy measurements of sample deformation. The applications include high-dimensional inverse problems with synthetic (Section 4.1) and experimental data (Section 4.2). Section 5 ends with conclusions and future directions.
@@ 2.2 Bayesian inference
$\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}\!\left({\mathbf{X}}={\mathbf{x}}|{\mathbf{Y}}=\hat{{\mathbf{y}}}\right)\propto\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}\!\left({\mathbf{Y}}=\hat{{\mathbf{y}}}|{\mathbf{X}}={\mathbf{x}}\right)\mathrm{p}_{{\mathbf{X}}}\!\left({{\mathbf{X}}={\mathbf{x}}}\right).$ (2)
@@ 2.3 Langevin dynamics and the score function
Markov chain Monte Carlo (MCMC) algorithms have been the staple of Bayesian machinery. MCMC methods are useful for sampling unnormalized probability models such as the posterior distribution in inverse problems. Early MCMC methods based on the Metropolis-Hastings (MH) algorithm [53, 54] use a proposal density to advance a Markov chain with an invariant distribution similar to the target posterior. However, in high-dimensional settings, simple MH-type MCMC methods are known to be inefficient as randomly chosen new states fail to explore high-likelihood regions quickly.
Langevin dynamics is an improvement over MH-MCMC, and uses a proposal density obtained from discretizing Langevin diffusion equation with a drift term informed by the gradient of the target posterior distribution. The resulting stochastic differential equation is defined as:
@@ 3 Bayesian inference using conditional score-based diffusion models
Bayesian inference methods try to approximate the posterior distribution of ${\mathbf{X}}$ for a given measurement $\hat{{\mathbf{y}}}$ using samples drawn from it. Eqs. 3 and 5 are useful for devising an MCMC method for sampling the posterior distribution. However, the ‘true’ score function remains inaccessible in practical applications. Using conditional score-based diffusion generative models, we approximate the score function using a neural network $\mathbf{s}({\mathbf{x}},{\mathbf{y}};\bm{\theta})$, popularly known as the score network, where $\bm{\theta}$ represents the parameters of the neural network. In this section, we develop the training objective for the score network, introduce the MCMC-based sampler called Annealed Langevin dynamics [31] that is useful for sampling target posteriors using a trained score network, provide guidelines regarding hyper-parameter selection and design choices for the score network, and finally provide an overview of the entire workflow.
@@ 3.1 Training the score network using denoising score matching
which is remarkable for one very particular reason. Note, estimating the objective $\mathcal{L}_{\mathrm{DSM}}(\bm{\theta})$ requires computing two expectations. The outer expectation is with respect to the joint distribution of ${\mathbf{X}}$ and ${\mathbf{Y}}$. We can easily sample the joint by first sampling ${\mathbf{X}}$ from the prior $\mathrm{p}_{{\mathbf{X}}}$ subsequently ${\mathbf{Y}}$ from the likelihood $\mathrm{p}_{{\mathbf{Y}}|{\mathbf{X}}}$, which simply requires the evaluation of the forward model Eq. 1. The inner expectation is with respect to the conditional distribution $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{X}}}$, which we can choose to be a tractable distribution that can be easily sampled. Thus, the Monte Carlo approximation of $\mathcal{L}_{\mathrm{DSM}}(\bm{\theta})$, which we denote as $\mathcal{L}_{\mathrm{eDSM}}(\bm{\theta})$, is given by
@@
where ${\mathbf{x}}^{(i)}$ and ${\mathbf{y}}^{(i)}$ are the ${i}^{\text{th}}$ realizations of ${\mathbf{x}}$ and ${\mathbf{y}}$ sampled from the joint $\mathrm{p}_{{\mathbf{X}}{\mathbf{Y}}}$. Moreover, ${\mathbf{y}}^{(i)}=\mathcal{F}({\mathbf{x}}^{(i)};\bm{\eta}^{(i)})$, where $\bm{\eta}^{(i)}$ is the ${i}^{\text{th}}$ realization of the noise vector.
@@ 3.2 Sampling using Annealed Langevin Dynamics
From Eq. 9, note, $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{Y}}}$ is the outcome of a convolution between the target posterior $\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}$ and the zero-mean Gaussian kernel with covariance matrix $\sigma^{2}\mathbb{I}_{{n_{\mathcal{X}}}}$. This is akin to a ’smoothing’ operation of the target posterior. In this work, we use Annealed Langevin Dynamics, proposed by Song and Ermon [31], to sample from the target posterior distributions using the learned score network starting from the smoothest target and gradually reducing the amount of smoothing; see Algorithm 1.
Input: Trained score network $\mathbf{s}({\mathbf{x}},{\mathbf{y}},\sigma;\bm{\theta})$, measurement $\hat{{\mathbf{y}}}$, sampling steps $T$, step size $\epsilon^{2}$, and perturbation kernel standard deviations $\left\{\sigma_{i}\right\}_{i=1}^{L}$
@@ 5             Sample ${\mathbf{z}}\sim\mathcal{N}(\bm{0},\mathbb{I}_{{n_{\mathcal{X}}}})$
6             ${\mathbf{x}}_{j}={\mathbf{x}}_{j-1}+\frac{\alpha_{i}}{2}\mathbf{s}({\mathbf{x}}_{j-1},\hat{{\mathbf{y}}},\sigma_{i};\bm{\theta})+\sqrt{\alpha_{i}}{\mathbf{z}}$
@@ 3.3 Hyper-parameters
There are a few hyper-parameters associated with the conditional score-based diffusion models. Some of these hyper-parameters are related to the training of the score network while others are associated with the sampling algorithm used to generate new realizations. Here we discuss how to choose the various important hyper-parameters.
The hyper-parameters associated with score estimation include the noise levels $\sigma_{1}$ through $\sigma_{L}$, and the number $L$ of noise levels. Following Song and Ermon [36], we choose $\sigma_{1}$ to be larger than the Euclidean distance between any two points of ${\mathbf{x}}$ within the training data. This helps smooth out the effect of $\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}$ from $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{Y}}}$ and sampling can commence from a near unimodal Gaussian distribution with covariance close to $\sigma_{1}^{2}\mathbb{I}_{{n_{\mathcal{X}}}}$. This helps avoid particles from getting stuck near modes of the target distribution and the sampling space can be properly explored with a limited number of particles. We choose $\sigma_{L}$ to be a very small value — typically $0.001$ or $0.01$ — so that the approximation $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{Y}}}(\tilde{{\mathbf{x}}}|{\mathbf{y}})\approx\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}({\mathbf{x}}|{\mathbf{y}})$ is better.
Following Song and Ermon [36], we choose $\sigma_{j}$ to be a geometric progression between $\sigma_{1}$ and $\sigma_{L}$, i.e., we fix $\gamma=\sigma_{j-1}/\sigma_{j}$. Moreover, we choose the number $L$ of noise levels such that particles from $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j-1}}|{\mathbf{Y}}}$ are good initializers for the Langevin dynamics MCMC when exploring $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j}}|{\mathbf{Y}}}$, which will be true when their is significant overlap between the two distribution. Song and Ermon [36] suggest choosing $\gamma$ such that:
@@
The hyper-parameters for the Annealed Langevin dynamics include the number of Langevin steps $T$ and the step size $\epsilon$. Increasing $L$ will increase the total number of Langevin steps, and the total time it takes to generate realizations from the target posterior distribution. So, we choose $T$ as large as possible as the compute budget allows. Finally, we choose $\epsilon$ such that particles from $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j-1}}|{\mathbf{X}}}$ can reach their steady state distribution $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j}}|{\mathbf{X}}}$ within $T$ steps. Song and Ermon [36] suggest choosing a value of $\epsilon$ for which [36]

### v3

# Conditional score-based diffusion models for solving inverse elasticity problems

@@ Abstract
journal: Computer Methods in Applied Mechanics and Engineering
@@ 1 Introduction
Clearly, conditional generative modeling resembles the process of solving an inverse problem within the Bayesian paradigm where ${\mathbf{X}}$ and ${\mathbf{Y}}$ now represent the quantity that must be inferred and the measurements available, respectively [17]. Thus, motivated by the recent wide-scale adoption of diffusion models, we proposed to use conditional diffusion models to solve a class of inverse problems in mechanics that involve inferring the spatially varying material constitutive parameters from full-field measurements of the mechanical response, such as displacements or strains. Inverse problems of this type have many applications in material characterization [18], structural health monitoring and damage detection [19, 20], and medical diagnosis [21, 22]. We are particularly motivated by the applications of this type of inverse problem in medical diagnosis, which arises when using elastography-based imaging techniques that assess the state of tissue using its mechanical properties, mainly to differentiate between healthy and diseased tissue [21, 23].
@@
Solving inverse problems of the aforementioned type can be challenging. First, the inferred quantity is a field that must be discretized to make the inverse problem tractable, leading to a high-dimensional inverse problem. Second, the forward model is often complex. Mechanistic models that simulate the deformations of the specimen in consideration may be a high-fidelity black-box model. Moreover, the measurement noise in these applications may be complex— non-Gaussian and non-additive. These factors combine to make Bayesian inference difficult since standard inference tools, such as Markov chain Monte Carlo (MCMC) methods and variational inference, may not perform well in the face of these challenges. However, the recent success of conditional diffusion models in solving high-dimensional inverse problems offers promise [24, 25, 26, 27, 28].
@@
We propose to use conditional score-based diffusion models (cSDMs) — a particular type of diffusion model that leverages deep neural networks to approximate the score function of the target posterior distribution [29, 30, 31]. In the context of inverse problems, the score function is the logarithm of the posterior density. The neural network approximating the score function is called the score network [31]. Approximating the score function allows greater flexibility in choosing the architecture for the score network in the absence of additional constraints necessary to model a valid probability density. For instance, the score network need not be invertible [31], unlike normalizing flows [32]. Moreover, a single neural network can approximate the posterior distribution for different realizations of the covariate ${\mathbf{Y}}$, which only requires adding a channel to the input of the neural network [33, 34, 35]. This also allows for re-using the trained score network for different realizations of ${\mathbf{Y}}$ during testing. Significantly, we show that the score network can be trained using a regression-type objective, see Eq. 15, which is much more stable than adversarial losses used to train GANs. Moreover, training only requires samples from the joint distribution between ${\mathbf{X}}$ and ${\mathbf{Y}}$. The training dataset can be easily created by simulating the forward model for different realizations of ${\mathbf{X}}$ sampled from a suitable prior or collected a priori through experiments. Thus, the forward model can be arbitrarily complex, black-box, and incompatible with automatic differentiation. This lends great utility to the proposed approach since most forward models for the inverse problems of practical interest are likely to have one or more of the aforementioned properties. Moreover, the training dataset is entirely synthetic and does not require access to experimental data, which can be scarce in some applications. After training the score network, we can sample the posterior distribution for any realization of ${\mathbf{Y}}$, perhaps the result of an experiment, using Langevin dynamics-based MCMC [36, 31, 37].
@@
We note that recent work by Jacobsen et al. [38] and Huang et al. [39] uses diffusion models for full-field inversion from sparse measurements. Jacobsen et al. [38] primarily focuses on enforcing physical consistency to the realizations generated using the conditional diffusion model, which is not the focus of the current work. Similarly, Huang et al. [39] train score-based diffusion models for unconditional generation on the joint space between the full-field coefficients and the corresponding solution. Measurement mismatch error and physics-informed residue terms help guide the sampling process to simultaneously recover missing observations and perform full-field inversion. The realizations obtained using cSDMs can be enhanced by enforcing physical consistency using ideas developed in [38, 39]. In contrast, Bastek et al. [40] propose to enforce physical consistency during training of the score network, which can also be used to augment the proposed approach. However, enforcing physical consistency by discretizing the underlying forward model may not be straightforward in many applications involving nonlinear phenomena and complex geometries. Therefore, we do not pursue informing the score network or the sampling process with physics knowledge in this work. Indeed, the proposed approach is agnostic to the physics of the forward process, i.e., we assume that the forward model is black-box.
@@
The remainder of this paper is organized as follows. In Section 2, we set up the inverse problem we intend to solve and provide the necessary background on score-based diffusion models. We introduce our proposed approach to solve inverse problems using cSDMs in Section 3 where Section 3.1 develops the training objective, Section 3.2 outlines the sampling scheme, Section 3.3 provides guidelines to choose various hyper-parameters associated with the model and sampling, Section 3.4 provides design details pertaining to the neural network used to approximate the score function, and Section 3.5 provides an overview of the various steps in the proposed approach. In Section 4, we use the proposed approach to solve various inverse problems that concern the inference of a spatially varying constitutive parameter from noisy measurements of sample deformation. The applications include high-dimensional inverse problems with synthetic (Section 4.1) and experimental data (Section 4.2). Section 5 ends with conclusions and future directions.
@@ 2.2 Bayesian inference
$\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}\!\left({\mathbf{X}}={\mathbf{x}}|{\mathbf{Y}}=\hat{{\mathbf{y}}}\right)\propto\mathrm{p}_{{\mathbf{Y}}|{\mathbf{X}}}\!\left({\mathbf{Y}}=\hat{{\mathbf{y}}}|{\mathbf{X}}={\mathbf{x}}\right)\mathrm{p}_{{\mathbf{X}}}\!\left({{\mathbf{X}}={\mathbf{x}}}\right).$ (2)
@@ 2.3 Langevin dynamics and the score function
MCMC algorithms have been the staple of Bayesian machinery. MCMC methods are useful for sampling unnormalized probability models such as the posterior distribution in inverse problems. Early MCMC methods based on the Metropolis-Hastings (MH) algorithm [55, 56] use a proposal density to advance a Markov chain with an invariant distribution similar to the target posterior. However, in high-dimensional settings, simple MH-type MCMC methods are known to be inefficient as randomly chosen new states fail to explore high-likelihood regions quickly.
Langevin dynamics is an improvement over MH-MCMC, and uses a proposal density obtained from discretizing Langevin diffusion equation with a drift term informed by the gradient of the target posterior distribution [36]. The resulting stochastic differential equation is defined as:
@@ 3 Bayesian inference using conditional score-based diffusion models
Bayesian inference methods try to approximate the posterior distribution of ${\mathbf{X}}$ for a given measurement $\hat{{\mathbf{y}}}$ using samples drawn from it. Eqs. 3 and 5 are useful for devising an MCMC method for sampling the posterior distribution. However, the ‘true’ score function remains inaccessible in practical applications. Using cSDMs, we approximate the score function using a neural network $\mathbf{s}({\mathbf{x}},{\mathbf{y}};\bm{\theta})$, popularly known as the score network, where $\bm{\theta}$ represents the parameters of the neural network. In this section, we develop the training objective for the score network, introduce the MCMC-based sampler called Annealed Langevin dynamics [31] that is useful for sampling target posteriors using a trained score network, provide guidelines regarding hyper-parameter selection and design choices for the score network, and finally provide an overview of the entire workflow.
@@ 3.1 Training the score network using denoising score matching
which is remarkable for one very particular reason. Note, estimating the objective $\mathcal{L}_{\mathrm{DSM}}(\bm{\theta})$ requires computing two expectations. The outer expectation is with respect to the joint distribution of ${\mathbf{X}}$ and ${\mathbf{Y}}$. We can easily sample the joint by first sampling ${\mathbf{X}}$ from the prior $\mathrm{p}_{{\mathbf{X}}}$ subsequently ${\mathbf{Y}}$ from the likelihood $\mathrm{p}_{{\mathbf{Y}}|{\mathbf{X}}}$, which simply requires the evaluation of the forward model Eq. 1. The inner expectation is with respect to the conditional distribution $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{X}}}$, which we can choose to be a tractable distribution that can be easily sampled. Thus, the Monte Carlo (MC) approximation of $\mathcal{L}_{\mathrm{DSM}}(\bm{\theta})$, which we denote as $\mathcal{L}_{\mathrm{eDSM}}(\bm{\theta})$, is given by
@@
where ${\mathbf{x}}^{(i)}$ and ${\mathbf{y}}^{(i)}$ are the ${i}^{\text{th}}$ realizations of ${\mathbf{x}}$ and ${\mathbf{y}}$, respectively, sampled from the joint $\mathrm{p}_{{\mathbf{X}}{\mathbf{Y}}}$. Moreover, ${\mathbf{y}}^{(i)}=\mathcal{F}({\mathbf{x}}^{(i)};\bm{\eta}^{(i)})$, where $\bm{\eta}^{(i)}$ is the ${i}^{\text{th}}$ realization of the noise vector.
@@ 3.2 Sampling using Annealed Langevin Dynamics
From Eq. 9, note, $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{Y}}}$ is the outcome of a convolution between the target posterior $\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}$ and the zero-mean Gaussian kernel with covariance matrix $\sigma^{2}\mathbb{I}_{{n_{\mathcal{X}}}}$. This is akin to a ‘smoothing’ operation of the target posterior. In this work, we use Annealed Langevin Dynamics, first proposed by Song and Ermon [31] and later improved in [37], to sample from the target posterior distributions using the learned score network starting from the smoothest target and gradually reducing the amount of smoothing; see Algorithm 1. Line 1 in Algorithm 1 is similar to Eq. 5 but uses the score network $\mathbf{s}$ to approximate the score function $\nabla_{{\mathbf{x}}}\log\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}$ at different levels of smoothing.
Input: Trained score network $\mathbf{s}({\mathbf{x}},{\mathbf{y}},\sigma;\bm{\theta})$, measurement $\hat{{\mathbf{y}}}$, sampling steps $T$, step size parameter $\epsilon$, and perturbation kernel standard deviations $\left\{\sigma_{i}\right\}_{i=1}^{L}$
@@ 5             Sample ${\mathbf{z}}\sim\mathcal{N}(\bm{0},\mathbb{I}_{{n_{\mathcal{X}}}})$
6             ${\mathbf{x}}_{j}={\mathbf{x}}_{j-1}+\alpha_{i}\mathbf{s}({\mathbf{x}}_{j-1},\hat{{\mathbf{y}}},\sigma_{i};\bm{\theta})+\sqrt{2\alpha_{i}}{\mathbf{z}}$
@@ 3.3 Hyper-parameters
There are a few hyper-parameters associated with cSDMs. Some of these hyper-parameters are related to the training of the score network while others are associated with the sampling algorithm used to generate new realizations. Here we discuss how to choose the various important hyper-parameters.
The hyper-parameters associated with score estimation include the noise levels $\sigma_{1}$ through $\sigma_{L}$, and the number $L$ of noise levels. Following Song and Ermon [37], we choose $\sigma_{1}$ to be larger than the Euclidean distance between any two points of ${\mathbf{x}}$ within the training data. This helps smooth out the effect of $\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}$ from $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{Y}}}$ and sampling can commence from a near unimodal Gaussian distribution with covariance close to $\sigma_{1}^{2}\mathbb{I}_{{n_{\mathcal{X}}}}$. This helps avoid realizations from getting stuck near modes of the target distribution and the sampling space can be properly explored with a limited number of realizations. We choose $\sigma_{L}$ to be a very small value — typically $0.001$ or $0.01$ — so that the approximation $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{Y}}}(\tilde{{\mathbf{x}}}|{\mathbf{y}})\approx\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}({\mathbf{x}}|{\mathbf{y}})$ is better. The approximation improves as $\sigma_{L}$ decreases because $\lim_{\sigma_{L}\to 0}\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{L}}|{\mathbf{Y}}}(\tilde{{\mathbf{x}}}_{\sigma_{L}}|{\mathbf{y}})=\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}(\tilde{{\mathbf{x}}}_{\sigma_{L}}|{\mathbf{y}})$ as $\mathrm{p}_{\tilde{{\mathbf{X}}}|{\mathbf{X}}}\!\left({\tilde{{\mathbf{x}}}|{\mathbf{x}}}\right)$ converges to the Dirac delta function with $\sigma_{L}\to 0$.
Following Song and Ermon [37], we choose $\sigma_{j}$ to be a geometric progression between $\sigma_{1}$ and $\sigma_{L}$, i.e., we fix $\gamma=\sigma_{j-1}/\sigma_{j}$. Moreover, we choose the number $L$ of noise levels such that realizations from $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j-1}}|{\mathbf{Y}}}$ are good initializers for the Langevin dynamics MCMC when exploring $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j}}|{\mathbf{Y}}}$, which will be true when there is significant overlap between the two distribution. Song and Ermon [37] suggest choosing $\gamma$ such that:
@@
The hyper-parameters for the Annealed Langevin dynamics include the number of Langevin steps $T$ and the step size parameter $\epsilon$. Increasing $L$ will increase the total number of Langevin steps, and the total time it takes to generate realizations from the target posterior distribution. So, we choose $T$ as large as possible as the compute budget allows. Finally, we choose $\epsilon$ such that realizations from $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j-1}}|{\mathbf{X}}}$ can reach their steady state distribution $\mathrm{p}_{\tilde{{\mathbf{X}}}_{\sigma_{j}}|{\mathbf{X}}}$ within $T$ steps. Song and Ermon [37] suggest choosing a value of $\epsilon$ for which [37]